# Splitting differential equations using Galois theory

### v1

@@ 4.3. Non-splitting when the binding group is an elliptic curve
We now want to use our methods to exhibit more uniformly internal maps that do not split. In [10, bottom of page 5], Jin and Moosa ask whether there exists a strongly minimal type $p\in S_{1}(F)$, with $F$ algebraically closed, such that $p$ is $\mathcal{C}$-internal with binding group an elliptic curve, with the type $\log_{\delta}^{-1}(p)$ almost $\mathcal{C}$-internal, but the map $\log_{\delta}:\log_{\delta}^{-1}(p)\rightarrow p$ not almost split. Using Lemma 3.15, we will show that there exist many such types.
To deal with algebraic closure issues, we will need to use the machinery of the $\mathrm{Ext}(\cdot,\cdot)$ functor classifying extensions of commutative algebraic groups. A good reference is [24, Chapter 7], which we will use in the rest of this section. We now recall some of the exposition and facts found in that chapter.
@@
{tikzcd}
commute. In that case, $f$ must be an isomorphism. For any two commutative algebraic groups $A$ and $B$, denote $\mathrm{Ext}(A,B)$ the set of isomorphism classes of commutative group extensions. We will now work with commutative groups and switch to additive notation. Here is a summary of the material of [24, VII.1] making $\mathrm{Ext}(\cdot,\cdot)$ into a functor.
@@
The following summarizes the properties we will need (see [24, Chapter 7, 1.1]):
Fact 4.12.
Let $\mathcal{A}$ be the category of commutative algebraic groups. The previous construction makes Ext(,)normal-$\mathrm{Ext}(\cdot,\cdot)$ into an additive bifunctor on $\mathcal{A}\times\mathcal{A}$, contravariant in the first coordinate and covariant in the second. In particular $\mathrm{Ext}(A,B)$ is always an algebraic group, and the maps $f_{*}$ and $g^{*}$ are morphisms of algebraic groups.
We will also need to identify the specific algebraic group $\mathrm{Ext}(E,G_{m})$, where $E$ is an elliptic curve. Recall that for any abelian variety $A$, we can form its dual abelian variety $\widetilde{A}$, which parametrizes the topologically trivial line bundles on $A$ (see [19, II.8]). The dual of an elliptic curve is also an elliptic curve. We then have the following (see [24, Chapter 7, 3.16]):
@@
The map $g:G_{m}\rightarrow G_{m}$ has finite kernel, denote it $H$. By [24, Chapter 7, 4.23, Theorem 12], the functor $\mathrm{Ext}(E,\cdot)$ is exact on the category of linear algebraic groups, and we thus obtain a short exact sequence:
@@
To do so, we show that there is a surjective morphism from $\operatorname{Hom}(\prescript{}{n}{E},H)$ to $\mathrm{Ext}(E,H)$, where $\prescript{}{n}{E}$ is the $n$-torsion of $E$. This torsion group is well-known to be finite (see [19, II.4]), so $\operatorname{Hom}(\prescript{}{n}{E},H)$ is finite as well.
Let $n\in\mathbb{N}$, we have maps $\nu_{n}:H\rightarrow H$ and $\mu_{n}:E\rightarrow E$ given by sending $x$ to $nx$. It is well known (see [19, II.4]) that $\mu_{n}$ is surjective and has finite kernel. Pick $n$ to be a multiple of the order of $H$, so that $\nu_{n}$ is the zero map, or in other words, so that $nh=0$ for all $h\in H$. We have a short exact sequence:
@@
By [24, Chapter 7, 1.2, Proposition 2], we obtain an exact sequence:
{tikzcd}
@@
To complete our proof, we will need to exhibit a $\mathcal{C}$-internal type with binding group isomorphic to a semiabelian variety. This follows from Kolchin’s solution to the inverse Galois problem for strongly normal extensions (see [14, Theorem 2] and also [15, Proposition 15.1]), as well as some translation to model-theoretic language. We give some details. Our exposition closely follows Marker’s in [17, Chapter 2, Section 9].
@@
Since $L/K$ is strongly normal, there is some tuple $\overline{a}$ such that $L=K\langle\overline{a}\rangle$. Let $p=\operatorname{tp}(\overline{a}/K)$. Marker’s proof of [17, Theorem 9.5] (and the discussion following it) shows that $\operatorname{tp}(\overline{a}/K)$ is $\mathcal{C}$-internal, and that there is an algebraic group $G$ such that $G(L/K)$ (resp. $\mathrm{Gal}(L/K)$) is isomorphic to $G(\mathcal{C}_{K})$ (resp. $G(\mathcal{C})$).
@@
Finally, we will use the Galois correspondence for binding groups. We refer the reader to [23, Theorem 2.3] for a proof, as well as the closest account to what we need that we could find. From that theorem, we deduce the following:
Fact 4.16.
@@
If $H$ is connected, we also obtain that $\operatorname{tp}(a/bF)$ is stationary (see [20, Chapter 1, Lemma 6.16]),.
@@
By Fact 4.13, there is an algebraic group $D$ that is a non-definably split extension of an elliptic curve $E$ by $G_{m}$, i.e. we have a map $g:D\rightarrow E$ with kernel $G_{m}$. Moreover, we can assume that $D$ does not belong to any finite subgroup of the elliptic curve $\mathrm{Ext}(E,G_{m})$ (equivalently $D$ is not a torsion point).
By Fact 4.15, there is an algebraically closed field $F$ and $p\in S(F)$ such that $p$ is $\mathcal{C}$-internal and weakly $\mathcal{C}$-orthogonal, with $\mathrm{Aut}_{F}(p/\mathcal{C})$ definably isomorphic to $D(\mathcal{C})$, where $G$ is a non-definably split extension of an elliptic curve by $G_{m}$. As was observed in [6, Fact 2.4], we may assume that $p$ is fundamental by taking a Morley power.
In particular the group $\mathrm{Aut}_{F}(p/\mathcal{C})$ has a definable Chevalley decomposition in the sense of [6, Fact 2.8], meaning a subgroup maximal among definable subgroups $F$-definably isomorphic to the $\mathcal{C}$-points of a linear algebraic group. This subgroup $L$ is called the linear part of $\mathrm{Aut}_{F}(p/\mathcal{C})$. It is easy to show, in this case, that the linear part of $\mathrm{Aut}_{F}(p/\mathcal{C})$ is the kernel of the map induced by $g$ on $\mathrm{Aut}_{F}(p/\mathcal{C})$ (in particular it is definably isomorphic to $G_{m}(\mathcal{C})$). Moreover, the linear part is $F$-definable and normal by [6, Fact 2.8].
By Fact 4.17, there is an $F$-definable fibration $f:p\rightarrow f(p)$ giving rise to the short exact sequence:
@@
Since both $p$ and $f(p)$ are fundamental, by [6, Lemma 2.7] $\mathrm{Aut}_{f(a)F}(p_{f(a)}/\mathcal{C})$ is definably isomorphic to $L$, and hence to $G_{m}(\mathcal{C})$. By Lemma 2.10 the binding groups $\mathrm{Aut}_{f(a)F}(p_{f(a)}/\mathcal{C})$ and $\mathrm{Aut}_{f(a)F}(\operatorname{tp}(b/f(a)F)/\mathcal{C})$ are isogenous, thus $\mathrm{Aut}_{f(a)\mathcal{C}}(\operatorname{tp}(b/f(a)F)/\mathcal{C})$ is isogenous to $G_{m}(\mathcal{C})$. Also $\mathrm{Aut}_{f(a)F}(\operatorname{tp}(b/f(a)F)/\mathcal{C})$ is a definable normal subgroup of $\mathrm{Aut}_{F}(s/\mathcal{C})$ by Fact 2.5.
As $L$ acts regularly on $p_{f(a)}$, we see that $U(a/f(a)F)=1$. We deduce from this that $U(s)=1$. As $s$ is $\mathcal{C}$-internal, this implies that $\mathrm{RM}(s)=1$, and as $F$ is algebraically closed, the type $s$ is strongly minimal. As $\mathrm{Aut}_{F}(s/\mathcal{C})$ has a normal subgroup isogenous to $G_{m}(\mathcal{C})$, Fact 4.3 implies that $\mathrm{Aut}_{F}(s/\mathcal{C})$ is definably isomorphic to $G_{m}(\mathcal{C})$. In fact we get that $\mathrm{Aut}_{f(a)F}(\operatorname{tp}(b/f(a)F)/\mathcal{C})=\mathrm{Aut}_{F}(s/\mathcal{C})$.
@@
is $F$-definably split by Lemmas 3.14 and 3.15.
@@
{tikzcd}
Both $\mathrm{Aut}_{F}(p/\mathcal{C})$ and $\mathrm{Aut}_{F}(f(p)\otimes s/\mathcal{C})$ are abelian, and therefore $F$-definably isomorphic to the $\mathcal{C}$-points of algebraic groups (see [6, Lemma 2.2] for a proof of that well-known fact). Denote the $\mathcal{C}$-points of these algebraic groups by $\overline{\mathrm{Aut}_{F}(p/\mathcal{C})}$ and $\overline{\mathrm{Aut}_{F}(f(p)\otimes s/\mathcal{C})}$. They are defined over $F\cap\mathcal{C}$, are isogenous, and by the same reasoning as in the previous paragraph, we obtain an $F\cap\mathcal{C}$-definable algebraic group $\overline{G}$ and a commutative diagram:
{tikzcd}
@@
By Fact 4.12, the maps $\iota$ and $\zeta$ give rise to two automorphisms $\iota_{*}$ and $\zeta_{*}$ of $\mathrm{Ext}(E,G_{m})$. We obtain that $\iota_{*}(\overline{\mathrm{Aut}_{F}(p/\mathcal{C})})=\overline{G}=\zeta_{*}(\overline{\mathrm{Aut}_{F}(s\otimes f(p)/\mathcal{C})})$.
We know that $\overline{\mathrm{Aut}_{F}(s\otimes f(p)/\mathcal{C})}$ is a $\mathcal{C}\cap F$-definably split extension of $E(\mathcal{C})$, therefore it is trivial as an element of $\mathrm{Ext}(E,G_{m})$, and so is its image $\overline{G}$, hence $\overline{G}$ is the identity of $\mathrm{Ext}(E,G_{m})$, and thus the extension $\overline{\mathrm{Aut}_{F}(p/\mathcal{C})}$ belongs to the kernel of $\iota_{*}$. By Proposition 4.14, the map $\iota_{*}$ has finite kernel. Therefore the extension $\overline{\mathrm{Aut}_{F}(p/\mathcal{C})}$, which is isomorphic to the extension $D$, belongs to a finite subgroup of $\mathrm{Ext}(E,G_{m})$. This contradicts our choice of $D$, as it was assumed that it did not belong to any finite subgroup of $\mathrm{Ext}(G_{m},E)=\widehat{E}$.
@@
There exists an algebraically closed field $F<\mathcal{U}$, a type $q\in S_{1}(F)$ such that $\log_{\delta}:q\rightarrow\log_{\delta}(q)$ does not almost-split, the type $\log_{\delta}(q)$ is strongly minimal and has a binding group isomorphic to the constant points of an elliptic curve $E$.
@@
Applying Lemma 4.7 to the algebraically closed field $F$ and $p\in S(F)$ obtained in Theorem 4.18, we obtain a type $q\in S_{1}(F)$, internal to $\mathcal{C}$, a fibration $\log_{\delta}:q\rightarrow\log_{\delta}(q)$ with $\log_{\delta}(q)\subset\operatorname{dcl}(f(p)(\mathcal{U})F)$. Because $f$ does not almost split, the map $\log_{\delta}:q\rightarrow\log_{\delta}(q)$ does not almost split either. To conclude, we just need to show that $\log_{\delta}(q)$ is strongly minimal and $\mathrm{Aut}_{F}(\log_{\delta}(q)/\mathcal{C})$ definably isomorphic to the constant points of an elliptic curve.
Because $\log_{\delta}(q)\subset\operatorname{dcl}(f(p)(\mathcal{U})F)$, we have that $\mathrm{Aut}_{F}(\log_{\delta}(q)/f(p),\mathcal{C})=\{\operatorname{id}\}$, therefore by Lemma 2.10 we obtain an $F$-definable surjective map $\mathrm{Aut}_{F}(f(p)/\mathcal{C})\rightarrow\mathrm{Aut}_{F}(\log_{\delta}(q)/\mathcal{C})$. As $\mathrm{Aut}_{F}(f(p)/\mathcal{C})$ is definably isomorphic to $E(\mathcal{C})$, and thus strongly minimal, its kernel is either the whole group, which would imply that $\log_{\delta}(q)$ is $\mathcal{C}$-definable (i.e. $\log_{\delta}(q)\subset\operatorname{dcl}(\mathcal{C},F)$), or finite.
Note that the type $\log_{\delta}(q)$ cannot be algebraic, as if it was, then $\log_{\delta}:q\rightarrow\log_{\delta}(q)$ would be almost split and thus so would $f$. Hence in the first possibility, we deduce that $\log_{\delta}(q)$ is not weakly orthogonal to $\mathcal{C}$. But any $\log_{\delta}(b)\models\log_{\delta}(q)$ is in the algebraic closure of some $a\models p$, which contradicts $p$ being weakly orthogonal to $\mathcal{C}$.
Therefore the kernel of the map is finite, and $\mathrm{Aut}_{F}(\log_{\delta}(q)/\mathcal{C})$ is definably isomorphic to an algebraic group isogenous to $E(\mathcal{C})$, which thus must be an elliptic curve. Note that by the proof of Lemma 4.7, for any $\log_{\delta}(b)\models q$, there is $f(a)\models f(p)$ such that $\log_{\delta}(b)\in\operatorname{dcl}(f(a),F)$. This, and the previous discussion, forces $\log_{\delta}(q)$ to be strongly minimal. ∎

### v3

@@ 4.3. Non-splitting when the binding group is an elliptic curve
We now want to use our methods to exhibit more uniformly internal maps that do not split. In [12, bottom of page 5], Jin and Moosa ask whether there exists a strongly minimal type $q\in S_{1}(F)$, with $F$ algebraically closed, such that $q$ is $\mathcal{C}$-internal with binding group an elliptic curve, with the type $\log_{\delta}^{-1}(q)$ almost $\mathcal{C}$-internal, but the map $\log_{\delta}:\log_{\delta}^{-1}(q)\rightarrow q$ not almost split. Using Lemma 3.12, we will show that there exist many such types.
To deal with algebraic closure issues, we will need to use the machinery of the $\mathrm{Ext}(\cdot,\cdot)$ functor classifying extensions of commutative algebraic groups. A good reference is [28, Chapter 7], which we will use in the rest of this section. We now recall some of the exposition and facts found in that chapter.
@@
[…]
commute. In that case, $f$ must be an isomorphism. For any two commutative algebraic groups $A$ and $B$, denote $\mathrm{Ext}(A,B)$ the set of isomorphism classes of commutative group extensions. We will now work with commutative groups and switch to additive notation. Here is a summary of the material of [28, VII.1] making $\mathrm{Ext}(\cdot,\cdot)$ into a functor.
@@
The following summarizes the properties we will need (see [28, Chapter 7, 1.1]):
@@
Let $\mathcal{A}$ be the category of commutative algebraic groups. The previous construction makes $\mathrm{Ext}(\cdot,\cdot)$ into an additive bifunctor on $\mathcal{A}\times\mathcal{A}$, contravariant in the first coordinate and covariant in the second. In particular $\mathrm{Ext}(A,B)$ is always an algebraic group, and the maps $f_{*}$ and $g^{*}$ are morphisms of algebraic groups.
We will also need to identify the specific algebraic group $\mathrm{Ext}(E,G_{m})$, where $E$ is an elliptic curve. Recall that for any abelian variety $A$, we can form its dual abelian variety $\widetilde{A}$, which parametrizes the topologically trivial line bundles on $A$ (see [21, II.8]). The dual of an elliptic curve is also an elliptic curve. We then have the following (see [28, Chapter 7, 3.16]):
Fact 4.14.
@@
The map $g:G_{m}\rightarrow G_{m}$ has finite kernel, denote it $H$. By [28, Chapter 7, 4.23, Theorem 12], the functor $\mathrm{Ext}(E,\cdot)$ is exact on the category of linear algebraic groups, and we thus obtain a short exact sequence:
@@
To do so, we show that there is a surjective morphism from $\operatorname{Hom}(\prescript{}{n}{E},H)$ to $\mathrm{Ext}(E,H)$, where $\prescript{}{n}{E}$ is the $n$-torsion of $E$. This torsion group is well-known to be finite (see [21, II.4]), so $\operatorname{Hom}(\prescript{}{n}{E},H)$ is finite as well.
Let $n\in\mathbb{N}$, we have maps $\nu_{n}:H\rightarrow H$ and $\mu_{n}:E\rightarrow E$ given by sending $x$ to $nx$. It is well known (see [21, II.4]) that $\mu_{n}$ is surjective and has finite kernel. Pick $n$ to be a multiple of the order of $H$, so that $\nu_{n}$ is the zero map, or in other words, so that $nh=0$ for all $h\in H$. We have a short exact sequence:
@@
By [28, Chapter 7, 1.2, Proposition 2], we obtain an exact sequence:
00{0}${\operatorname{Hom}(E,H)}$${\operatorname{Hom}(E,H)}$${\operatorname{Hom}(\prescript{}{n}{E},H)}$${\mathrm{Ext}(E,H)}$${\mathrm{Ext}(E,H)}$Ext(${\mathrm{Ext}(\prescript{}{n}{E},H)}$$\scriptstyle{\cdot\circ\mu_{n}}$$\scriptstyle{\cdot\circ\iota}$$\scriptstyle{d}$$\scriptstyle{\mu_{n}^{*}}$$\scriptstyle{\iota^{*}}$
@@
To complete our proof, we will need to exhibit a $\mathcal{C}$-internal type with binding group isomorphic to a semiabelian variety. This follows from Kolchin’s solution to the inverse Galois problem for strongly normal extensions (see [15, Theorem 2] and also [17, Proposition 15.1]), as well as some translation to model-theoretic language. We give some details. Our exposition closely follows Marker’s in [19, Chapter 2, Section 9].
@@
Since $L/K$ is strongly normal, there is some tuple $\overline{a}$ such that $L=K\langle\overline{a}\rangle$. Let $p=\operatorname{tp}(\overline{a}/K)$. Marker’s proof of [19, Theorem 9.5] (and the discussion following it) shows that $\operatorname{tp}(\overline{a}/K)$ is $\mathcal{C}$-internal, and that there is an algebraic group $G$ such that $G(L/K)$ (resp. $\mathrm{Gal}(L/K)$) is isomorphic to $G(\mathcal{C}_{K})$ (resp. $G(\mathcal{C})$).
@@
Finally, we will use the Galois correspondence for binding groups. We refer the reader to [27, Theorem 2.3] for a proof, as well as the closest account to what we need that we could find. From that theorem, we deduce the following:
Fact 4.17.
@@
If $H$ is connected, we also obtain that $\operatorname{tp}(a/bF)$ is stationary (see [22, Chapter 1, Lemma 6.16]).
@@
By Fact 4.14, there is an algebraic group $D$ that is a non-definably split extension of an elliptic curve $E$ by $G_{m}$, i.e. we have a map $g:D\rightarrow E$ with kernel $G_{m}$. Moreover, we can assume that $D$ does not belong to any finite subgroup of the elliptic curve $\mathrm{Ext}(E,G_{m})$ (equivalently $D$ is not a torsion point).
By Fact 4.16, there is an algebraically closed field $F$ and $p\in S(F)$ such that $p$ is $\mathcal{C}$-internal and weakly $\mathcal{C}$-orthogonal, with $\mathrm{Aut}_{F}(p/\mathcal{C})$ definably isomorphic to $D(\mathcal{C})$, where $G$ is a non-definably split extension of an elliptic curve by $G_{m}$. As was observed in [8, Fact 2.4], we may assume that $p$ is fundamental by taking a Morley power.
In particular the group $\mathrm{Aut}_{F}(p/\mathcal{C})$ has a definable Chevalley decomposition in the sense of [8, Fact 2.8], meaning a subgroup maximal among definable subgroups $F$-definably isomorphic to the $\mathcal{C}$-points of a linear algebraic group. This subgroup $L$ is called the linear part of $\mathrm{Aut}_{F}(p/\mathcal{C})$. It is easy to show, in this case, that the linear part of $\mathrm{Aut}_{F}(p/\mathcal{C})$ is the kernel of the map induced by $g$ on $\mathrm{Aut}_{F}(p/\mathcal{C})$ (in particular it is definably isomorphic to $G_{m}(\mathcal{C})$). Moreover, the linear part is $F$-definable and normal by [8, Fact 2.8].
By Fact 4.18, there is an $F$-definable fibration $f:p\rightarrow f(p)$ giving rise to the short exact sequence:
@@
Since both $p$ and $f(p)$ are fundamental, by [8, Lemma 2.7] $\mathrm{Aut}_{f(a)F}(p_{f(a)}/\mathcal{C})$ is definably isomorphic to $L$, and hence to $G_{m}(\mathcal{C})$. By Lemma 2.10 the binding groups $\mathrm{Aut}_{f(a)F}(p_{f(a)}/\mathcal{C})$ and $\mathrm{Aut}_{f(a)F}(\operatorname{tp}(b/f(a)F)/\mathcal{C})$ are isogenous, thus $\mathrm{Aut}_{f(a)\mathcal{C}}(\operatorname{tp}(b/f(a)F)/\mathcal{C})$ is isogenous to $G_{m}(\mathcal{C})$. Also $\mathrm{Aut}_{f(a)F}(\operatorname{tp}(b/f(a)F)/\mathcal{C})$ is a definable subgroup of $\mathrm{Aut}_{F}(s/\mathcal{C})$ by Fact 2.5.
As $L$ acts regularly on $p_{f(a)}$, we see that $U(a/f(a)F)=1$. We deduce from this that $U(s)=1$. As $s$ is $\mathcal{C}$-internal, this implies that $\mathrm{RM}(s)=1$, and as $F$ is algebraically closed, the type $s$ is strongly minimal.
Fact 4.2 rules out the cases of the additive group and an elliptic curve for $\mathrm{Aut}_{F}(s/\mathcal{C})$. By Lemma 2.10, the groups $\mathrm{Aut}_{F}(p/\mathcal{C})$ and $\mathrm{Aut}_{F}(s\otimes f(p)/\mathcal{C})$ are isogenous. The projection map $s\otimes f(p)\rightarrow s$ yields a definable surjective morphism $\mathrm{Aut}_{F}(s\otimes f(p)/\mathcal{C})\rightarrow\mathrm{Aut}_{F}(s/\mathcal{C})$. As the former is isogenous to a semi-abelian variety, it cannot have any definable morphism to $G_{a}\ltimes G_{m}$ nor $\mathrm{PSL_{2}}$, ruling out the other two cases. Thus $\mathrm{Aut}_{F}(s/\mathcal{C})$ is definably isomorphic to $G_{m}(\mathcal{C})$, and by connectedness, we get $\mathrm{Aut}_{f(a)F}(\operatorname{tp}(b/f(a)F)/\mathcal{C})=\mathrm{Aut}_{F}(s/\mathcal{C})$.
@@
is $F$-definably split by Lemmas 3.11 and 3.12.
@@
[…]
Both $\mathrm{Aut}_{F}(p/\mathcal{C})$ and $\mathrm{Aut}_{F}(f(p)\otimes s/\mathcal{C})$ are abelian, and therefore $F$-definably isomorphic to the $\mathcal{C}$-points of algebraic groups (see [8, Lemma 2.2] for a proof of that well-known fact). Denote the $\mathcal{C}$-points of these algebraic groups by $\overline{\mathrm{Aut}_{F}(p/\mathcal{C})}$ and $\overline{\mathrm{Aut}_{F}(f(p)\otimes s/\mathcal{C})}$. They are defined over $F\cap\mathcal{C}$, are isogenous, and by the same reasoning as in the previous paragraph, we obtain an $F\cap\mathcal{C}$-definable algebraic group $\overline{G}$ and a commutative diagram:
[…]
@@
By Fact 4.13, the maps $\iota$ and $\zeta$ give rise to two automorphisms $\iota_{*}$ and $\zeta_{*}$ of $\mathrm{Ext}(E,G_{m})$. We obtain that $\iota_{*}(\overline{\mathrm{Aut}_{F}(p/\mathcal{C})})=\overline{G}=\zeta_{*}(\overline{\mathrm{Aut}_{F}(s\otimes f(p)/\mathcal{C})})$.
We know that $\overline{\mathrm{Aut}_{F}(s\otimes f(p)/\mathcal{C})}$ is a $\mathcal{C}\cap F$-definably split extension of $E(\mathcal{C})$, therefore it is trivial as an element of $\mathrm{Ext}(E,G_{m})$, and so is its image $\overline{G}$, hence $\overline{G}$ is the identity of $\mathrm{Ext}(E,G_{m})$, and thus the extension $\overline{\mathrm{Aut}_{F}(p/\mathcal{C})}$ belongs to the kernel of $\iota_{*}$. By Proposition 4.15, the map $\iota_{*}$ has finite kernel. Therefore the extension $\overline{\mathrm{Aut}_{F}(p/\mathcal{C})}$, which is isomorphic to the extension $D$, belongs to a finite subgroup of $\mathrm{Ext}(E,G_{m})$. This contradicts our choice of $D$, as it was assumed that it did not belong to any finite subgroup of $\mathrm{Ext}(G_{m},E)=\widehat{E}$. ∎
@@
There exists an algebraically closed field $F<\mathcal{U}$, a $\mathcal{C}$-internal type $p\in S_{1}(F)$ such that $\log_{\delta}:p\rightarrow\log_{\delta}(p)$ does not almost-split, the type $\log_{\delta}(p)$ is strongly minimal and has a binding group isomorphic to the constant points of an elliptic curve $E$.
@@
Applying Lemma 4.3 to the algebraically closed field $F$ and $p\in S(F)$ obtained in Theorem 4.19, we obtain a type $p^{\prime}\in S_{1}(F)$, internal to $\mathcal{C}$, a fibration $\log_{\delta}:p^{\prime}\rightarrow\log_{\delta}(p^{\prime})$ with $\log_{\delta}(p^{\prime})\subset\operatorname{dcl}(f(p)(\mathcal{U})F)$. Because $f$ does not almost split, the map $\log_{\delta}:p^{\prime}\rightarrow\log_{\delta}(p^{\prime})$ does not almost split either. To conclude, we just need to show that $\log_{\delta}(p^{\prime})$ is strongly minimal and $\mathrm{Aut}_{F}(\log_{\delta}(p^{\prime})/\mathcal{C})$ definably isomorphic to the constant points of an elliptic curve.
Because $\log_{\delta}(p^{\prime})\subset\operatorname{dcl}(f(p)(\mathcal{U})F)$, we have that $\mathrm{Aut}_{F}(\log_{\delta}(p^{\prime})/f(p),\mathcal{C})=\{\operatorname{id}\}$, therefore by Lemma 2.10 we obtain an $F$-definable surjective map $\mathrm{Aut}_{F}(f(p)/\mathcal{C})\rightarrow\mathrm{Aut}_{F}(\log_{\delta}(p^{\prime})/\mathcal{C})$. As $\mathrm{Aut}_{F}(f(p)/\mathcal{C})$ is definably isomorphic to $E(\mathcal{C})$, and thus strongly minimal, its kernel is either the whole group, which would imply that $\log_{\delta}(p^{\prime})$ is $\mathcal{C}$-definable (i.e. $\log_{\delta}(p^{\prime})\subset\operatorname{dcl}(\mathcal{C},F)$), or finite.
Note that the type $\log_{\delta}(p^{\prime})$ cannot be algebraic, as if it was, then $\log_{\delta}:p^{\prime}\rightarrow\log_{\delta}(p^{\prime})$ would be almost split and thus so would $f$. Hence in the first possibility, we deduce that $\log_{\delta}(p^{\prime})$ is not weakly orthogonal to $\mathcal{C}$. But any $\log_{\delta}(b)\models\log_{\delta}(p^{\prime})$ is in the algebraic closure of some $a\models p$, which contradicts $p$ being weakly orthogonal to $\mathcal{C}$.
Therefore the kernel of the map is finite, and $\mathrm{Aut}_{F}(\log_{\delta}(p^{\prime})/\mathcal{C})$ is definably isomorphic to an algebraic group isogenous to $E(\mathcal{C})$, which thus must be an elliptic curve. Note that by the proof of Lemma 4.3, for any $\log_{\delta}(b)\models p^{\prime}$, there is $f(a)\models f(p)$ such that $\log_{\delta}(b)\in\operatorname{dcl}(f(a),F)$. This, and the previous discussion, forces $\log_{\delta}(p^{\prime})$ to be strongly minimal. ∎